OpenStudy (anonymous):

(4^16) ^1/2 How can I simplify this expression

OpenStudy (anonymous):

You can perform a few operations on exponents. \[a^x * a^y = a^{x+y}\]\[\frac{ a^x }{ a^y }=a^{x-y}\]\[(a^x)^y=a^{x*y}\]The last one is what you need. Substitute in 16 for x and 1/2 for y, then solve. Should be pretty easy from there.